# Multi-modal Data Spectrum: Multi-modal Datasets are Multi-dimensional

### v1

@@ 4.3 Category analysis
Our analysis shows that aggregate performance metrics on multi-modal benchmarks can be misleading, as they obscure strong uni-modal biases at the sub-category level. As demonstrated in Figure 4, benchmarks that ostensibly rely on inter-modality dependencies (Figure 2) contain significant intra-modality dependencies when evaluated at a more granular level.
@@
Figure 4: Analysis of sub-categories across datasets showing dependency on individual modalities. Although benchmarks may be designed for inter-modality reasoning, we show a strong dependence on text for categories such as relative location in ADE and $CO$, or higher-grade questions in ScienceQA and multiple categories in MMMU and MMMUPro. This highlights how aggregate metrics can obscure that many instances may not require multi-modal reasoning. We show standard accuracy in yellow and contributions from text in blue, image in green, and random in orange.
This discrepancy is evident across several datasets. In ADE (Zhou et al., 2019) and COCO (Lin et al., 2014; Tong et al., 2024a), while a text-only model’s overall performance is only marginally above chance (see Figure 2), it achieves substantial accuracy on the relative location sub-category (Figures 4(a) and 4(b)). This phenomenon is amplified in knowledge-intensive benchmarks. In ScienceQA (Lu et al., 2022) (Figure 4(d)), text-only performance accounts for the majority of the accuracy of questions aimed at grades 10-12. Likewise, many academic subjects within the MMMU and MMMU Pro benchmarks (Yue et al., 2024) (Figures 4(e) and 4(f)) contain many instances solvable with a question or an image, respectively, allowing unimodal models to succeed without question or visual information. Conversely, Q-Bench (Wu et al., 2023) (Figure 4(c)) exhibits the opposite pattern. Individual categories show a dependence on both image and text intra-modality dependencies, yet the aggregate metrics in Figure 2 indicate a notable bias toward the image modality.
These findings are further corroborated by our analysis of datasets such as MME (Fu et al., 2023) and BLINK (Fu et al., 2024) in Figure A.6. We demonstrate that the degree of modality dependence is often inconsistent within a single benchmark. This highlights the profound difficulty in curating well-balanced multi-modal datasets, as uni-modal dependencies can emerge and vary unpredictably across different sub-populations of the data.
@@ 5 Limitations and Future Work
To more holistically evaluate multi-modal capabilities, we propose two crucial future directions. First, we should progress towards open-ended answer generation and evaluation (Rei et al., 2020; Balepur et al., 2025). Evaluating free-form responses presents significant challenges. The same meaning can be expressed in many ways, making automated evaluation difficult. This often requires human evaluation, which is slow and expensive. Therefore, we believe advancing these evaluation methods is a critical goal for reliably measuring multi-modal capabilities.
Second, models must be equipped with the ability to abstain from answering when presented with ambiguous or irrelevant inputs (Whitehead et al., 2022; Feng et al., 2024; Stengel-Eskin et al., 2024). We conduct a preliminary experiment with OpenAI GPT-5 and Google Gemini 2.5 Pro, showing cases where the image or the question was irrelevant to the answer in Figure 5. Despite facilitating abstention by augmenting the instruction set with a “None of the above” option, this approach is largely insufficient to overcome the dependence on uni-modal dependencies for both GPT-5 and Gemini 2.5 Pro models. This highlights that models have a tendency to generate a plausible-sounding but incorrect response over acknowledging ambiguity or lack of information with confidence. An important direction for future research is to develop methods for meaningful abstention.
@@ 6 Conclusion
Our work dissects the uni-modal and multi-modal dependencies of MLLMs on 24 benchmarks. Our analysis reveals that the degree to which each benchmark exhibits these dependencies varies significantly, not only across different datasets but also within them. We found that efforts to mitigate text-based dependencies have often unintentionally resulted in new image-based, perpetuating a cycle of superficial fixes. This suggests that progress cannot be achieved simply by developing more benchmarks or chasing leaderboard metrics. Instead, we must critically assess existing evaluation methods, move beyond standard multiple-choice formats and train models to abstain when an answer cannot be confidently determined. Meaningful advances in multimodal learning require understanding how a model arrives at an answer, not just what answer it provides.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research, Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI), National Science Foundation (NSF) award No. 1922658, Center for Advanced Imaging Innovation and Research (CAI2R), National Center for Biomedical Imaging and Bioengineering operated by NYU Langone Health, and National Institute of Biomedical Imaging and Bioengineering through award number P41EB017183.
Appendix A Supplementary Material
Organization.
In the supplementary material, we provide the implementation details in Section A.1 and additional results in Section A.2.
A.1 Experimental Details
@@
We use the Cambrian-1 [Tong et al., 2024a] open-sourced codebase for all the experiments. We use their publicly released models for evaluation. Datasets like AI2D [Kembhavi et al., 2016], ChartQA [Masry et al., 2022], MMBench [Liu et al., 2024a], MME [Fu et al., 2023], MMMU [Yue et al., 2024], MMVet, POPE [Li et al., 2023b], RealWorldQA [xAI, 2024], SEED [Li et al., 2023a], TextVQA [Singh et al., 2019], and VizWiz [Gurari et al., 2018] were sourced from LMMS-eval, while others such as ADE, Blink [Fu et al., 2024], COCO [Lin et al., 2014], GQA [Hudson and Manning, 2019], MathVista [Lu et al., 2023], MMMUPro [Yue et al., 2024], MMStar [Chen et al., 2024], MMVP [Tong et al., 2024b], OCRBench [Liu et al., 2024b], Omni3D [Brazil et al., 2023, Tong et al., 2024a], QBench [Wu et al., 2024], ScienceQA [Lu et al., 2022], and $V^{*}$Bench [Wu and Xie, 2024] were used from their respective sources. We did not use the datasets that required external submissions such as DocVQA and InfoVQA.
A.2 Additional Results
We show the effect of model size on additional datasets in Figure A.7 and results on categories for MME and BLINK in Figure A.6.
@@ A.2 Additional Results
Figure A.6: Analysis of sub-categories for MME and BLINK dataset.
@@
(b) MMStar
(c) MMBench (en)
(d) SEED-Bench
(e) SEED-Bench
(f) MMMUPro
(g) GQA
(h) TextVQA
Figure A.7: Effect of Model Size on Additional Datasets. Performance of various models (8B, 13B, 34B, and a majority-vote ensemble).

### v2

@@ 4.3 Category analysis
Aggregate performance metrics can be misleading, often obscuring strong intra-modality dependencies at the sub-category level. As shown in Figure 5, the benchmarks that appeared to use inter-modality dependencies in Figure 2 also contain intra-modality dependencies when evaluated at a granular level.
@@
Figure 5: Analysis of sub-categories across datasets showing dependency on individual modalities. Although benchmarks may be designed for inter-modality reasoning, we show a strong dependence on text for categories such as relative location in ADE and COCO, or higher-grade questions in ScienceQA and multiple categories in MMMU and MMMUPro. This highlights how aggregate metrics can obscure that many instances may not require multi-modal reasoning. We show standard accuracy in yellow and contributions from text in blue, image in green, and random in orange.
This discrepancy is evident across several datasets. In ADE (Zhou et al., 2019) and COCO (Lin et al., 2014; Tong et al., 2024a), while a text-only model’s overall performance is only marginally above chance (see Figure 2), it achieves substantial accuracy on the relative location sub-category (Figures 5(a) and 5(b)). This phenomenon is amplified in knowledge-intensive benchmarks. In ScienceQA (Lu et al., 2022) (Figure 5(d)), text-only performance accounts for the majority of the accuracy of questions aimed at grades 10-12. Likewise, many academic subjects within the MMMU and MMMU Pro benchmarks (Yue et al., 2024) (Figures 5(e) and 5(f)) contain many instances solvable with a question or an image, respectively, allowing unimodal models to succeed without question or visual information. Conversely, Q-Bench (Wu et al., 2023) (Figure 5(c)) exhibits the opposite pattern. Individual categories show a dependence on both image and text intra-modality dependencies, yet the aggregate metrics in Figure 2 indicate a notable bias toward the image modality.
These findings are corroborated by our analysis of datasets such as MME (Fu et al., 2023) and BLINK (Fu et al., 2024) in Figure B.7. We demonstrate that the degree of modality dependence is often inconsistent within a single benchmark. This highlights the multi-dimensional nature of multi-modal datasets, intra- and inter-modality dependencies emerge and vary unpredictably across different sub-populations of the data.
@@ 5 Limitations and Future Work
To more holistically evaluate multi-modal capabilities, we propose several crucial future directions. First, we should progress towards building benchmarks that focus on open-ended answer generation and evaluation (Rei et al., 2020; Balepur et al., 2025). Evaluating free-form responses presents significant challenges. The same meaning can be expressed in many ways, making automated evaluation difficult. This often requires human evaluation, which is slow and expensive. We believe progress in this direction is essential for measuring the necessary multi-modal capabilities.
Second, both benchmarks and models must support the ability to abstain from answering when presented with ambiguous inputs (Whitehead et al., 2022; Feng et al., 2024; Stengel-Eskin et al., 2024). We conduct a preliminary experiment with OpenAI GPT-5 and Google Gemini 2.5 Pro, showing cases where the image or the question was irrelevant to the answer in Figure 6. Despite facilitating abstention by augmenting the instruction set with a “None of the above” option, this approach is largely insufficient to overcome the dependence on uni-modal dependencies for both the models. This highlights that models have a tendency to generate a plausible-sounding but incorrect response over acknowledging ambiguity or lack of information with confidence. Future work should prioritize methods to encourage meaningful abstention.
Lastly, we encourage future benchmarks and models to report not only aggregated performance but also modality-specific and random baselines to better measure progress. From a benchmark perspective, this helps in understanding how a new dataset compares with existing ones in the inherent biases. From a model perspective, it clarifies where performance gains come from and guides meaningful future improvements.
@@ 6 Conclusion
Our work critically dissects the intra- and inter-modality dependencies of MLLMs on 23 benchmarks. We show that no dataset is truly multi-modal, as each measures multiple dimensions of multi-modal learning with different strengths of intra- and inter-modality dependence. These strengths vary substantially both across benchmarks and across categories within the same benchmark. We find that efforts to mitigate text-based dependencies have often introduced new image-based dependencies, perpetuating a cycle of superficial fixes. This suggests that meaningful progress cannot be achieved simply by developing more benchmarks or chasing leaderboard metrics. Instead, we must critically assess the existing evaluation methods. This includes moving beyond standard multiple-choice formats, incorporating scenarios where models should abstain when they are uncertain, and examining how a model arrives at an answer rather than only what answer it produces.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research, Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI), National Science Foundation (NSF) award No. 1922658, Center for Advanced Imaging Innovation and Research (CAI2R), National Center for Biomedical Imaging and Bioengineering operated by NYU Langone Health, and National Institute of Biomedical Imaging and Bioengineering through award number P41EB017183. The computational requirements for this work were supported by NYU IT High Performance Computing resources, services, and staff expertise and NYU Langone High Performance Computing Core’s resources and personnel. This work was partly supported in part by the NYUAD Center for Interdisciplinary Data Science & AI (CIDSAI), funded by Tamkeen under the NYUAD Research Institute Award CG016. This content is solely the responsibility of the authors and does not represent the views of the funding agencies.
Appendix
Organization
We provide the implementation details in Appendix A and additional results in Appendix B.
Appendix A Experimental Details
@@
We use the Cambrian-1 [Tong et al., 2024a] open-sourced codebase for all the experiments. We use their publicly released models for evaluation. Datasets like AI2D [Kembhavi et al., 2016], ChartQA [Masry et al., 2022], MMBench [Liu et al., 2024b], MME [Fu et al., 2023], MMMU [Yue et al., 2024], POPE [Li et al., 2023b], RealWorldQA [xAI, 2024], SEED [Li et al., 2023a], TextVQA [Singh et al., 2019], and VizWiz [Gurari et al., 2018] were sourced from LMMS-eval, while others such as ADE, Blink [Fu et al., 2024], COCO [Lin et al., 2014], GQA [Hudson and Manning, 2019], MathVista [Lu et al., 2023], MMMUPro [Yue et al., 2024], MMStar [Chen et al., 2024], MMVP [Tong et al., 2024b], OCRBench [Liu et al., 2024c], Omni3D [Brazil et al., 2023, Tong et al., 2024a], Q-Bench [Wu et al., 2024], ScienceQA [Lu et al., 2022], and $V^{*}$Bench [Wu and Xie, 2024] were used from their respective sources.
Appendix B Additional Results
The results on different categories for MME [Fu et al., 2023] and BLINK [Fu et al., 2024] datasets are shown in Figure B.7. We further show the effect of model size on additional datasets in Figure B.8, Figure B.9, Figure B.10 and the effect of model types on additional datasets in Figure B.11.
@@ Appendix B Additional Results
Figure B.7: Analysis of sub-categories for MME and BLINK dataset.
(a) TextVQA
(b) ChartQA
(c) COCO
(d) MME
Figure B.8: Effect of model size on additional datasets. Performance of various models (8B, 13B, 34B, and a majority-vote ensemble) with both image and text intra-modality dependencies (top) and primarily inter-modality dependency (bottom). The bars represent standard accuracy and contributions from text, image, and random (bars are in the same order).
(a) MMStar
(b) MMBench (en)
(c) SEED-Bench
(d) Q-Bench
Figure B.9: Effect of model size on additional datasets with image intra-modality dependency. Performance of various models (8B, 13B, 34B, and a majority-vote ensemble). The bars represent standard accuracy and contributions from text, image, and random (bars are in the same order).
@@
(b) MMMU Pro
(c) GQA
(d) ScienceQA
(e) Omni
(f) MathVista
Figure B.10: Effect of model size on additional datasets with text intra-modality dependency. Performance of various models (8B, 13B, 34B, and a majority-vote ensemble). The bars represent standard accuracy and contributions from text, image, and random (bars are in the same order).
(a) MMStar
(b) MMMU
(c) OCRBench
(d) Omni
(e) QBench
(f) ScienceQA
(g) TextVQA
(h) $V^{*}$ Bench
Figure B.11: Effect of model types on additional datasets. Performance comparison between various models such as LLava-Next (May 2024), Cambrian-1 8b (June 2024), Qwen2.5-VL (April 2025) and Qwen3-VL (October 2025). The bars represent standard accuracy and attributed contributions from text, image, and random (bars are in the same order).